OpenStudy (hockeychick23):

Suppose the probability of success in a binomial event is 0.15. What is the probability of failure?

OpenStudy (anonymous):

P(failure) = 1 - P(success)

OpenStudy (hockeychick23):

ok so P(failure)= 1- 0.15 so failure= 0.85?

OpenStudy (anonymous):

Correct. :)

OpenStudy (hockeychick23):

thanks!

OpenStudy (anonymous):

anytime :)